solve for f o g and g o f .............f(x)=5sqrt(2x+1) and g(x)=(1/2)x^5-(1/2)
Answer by stanbon(48568)

solve for f o g and g o f .............f(x)=5sqrt(2x+1) and g(x)=(1/2)x^5-(1/2)
----
fog=f[(1/2)x^5-(1/2)]
= 5sqrt[2[1/2x^5-(1/2)]]
=5[sqrt(x^5-1)]
gof=g[5sqrt(2x+1)]
= [(1/2)(5sqrt(2x+1))^5-(1/2)]
= (5/2)(2x+1)^2(2x+1)-(1/2)
==========
Cheers,
Stan H.